A swimmer coming out of the pool is covered with a film of water about 80gram. .how much heat must supplied to evaporate water?

Grade:12

1 Answers

Vikas TU
7 years ago
80 grams means
moles of water = 80/18 = 40/9 =4.44 moles.
W = nRdeltaT = 4.44*R*(373 – T).
Now here U havent provided the full information.
U must provided the room temp.
and also enthalpy change or internal enrgy.